A will states that john is ot get 3 times as much money as mary. the total amount they will receive is $11,000.

a.write a system of equations describing this situation
b.solve to find the amounts of money john and Mary will get.

Answer by dfrazzetto(283)


J + M = 11000
J = 3M


----------------


3M + M = 11000
4M = 11000
M = $2750
J = 3(2750) = $8250